What are the roots of the polynomial equation? Need help lease asap to graduate!!
(2x -1)(x - 8)(x+1) = 0
Select each correct answer?
-8
-1
-1/2
0
1/2
1
8

Answer by Boreal(15235)

Set each one equal to 0 and solve for x
(2x-1)=0
x=1/2
x-8=0
x=8
x+1=0
x=-1
(1/2, 8, -1)